40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Nov 30 12:02:05 2017, awade, Craig, DailyProgress, scatter, Initial buzz test and cleaning of periscope mirrors and final MM lens 20171129_135353noiseBudget.pdf2017-11-29_21.30.17.jpg2017-11-29_21.32.12.jpg
    Reply  Tue Dec 19 22:44:48 2017, awade, DailyProgress, scatter, Buzz test of beat board 20171119_BeatBoard_BuzzMap.JPG
    Reply  Thu Dec 21 18:57:12 2017, awade, DailyProgress, scatter, Switch out final MM lens in North path IMG_1464.JPG20171121_ReplacedFinalNorthMMLens.jpg
       Reply  Fri Dec 22 14:38:24 2017, awade, DailyProgress, scatter, Fixed bad PDH error signal 
Message ID: 1994     Entry time: Thu Nov 30 12:02:05 2017     Reply to this: 2022   2026
Author: awade, Craig 
Type: DailyProgress 
Category: scatter 
Subject: Initial buzz test and cleaning of periscope mirrors and final MM lens 

--- This is a backpost from last week ---

The FSS and PLL are now pretty much locked all the time with the last reported settings (see PLL:1991). The PLL and FSS have all remained in locked state for periods of over 48 hours.

Craig and I had a first look at the scatter noise shelf with a white noise PZT buzzer to see where where the worst offending sources of parasitic interferometers would be.  The periscopes and around the PDs seem like the most likely points from our initial tests.

There was a lot of foam dust on the upward facing periscope mirrors (see pictures PSL:1980) and on the final MM mirrors before the ref cavs.  We gave these a clean as an initial test.  The north path 77.3 mm ROC mirror directly before the periscope has a permanent burn mark in the center: see picture below.  I tried a gentle wipe with methanol and then a little more of an aggressive scrub and it wont budge. Craig has cleaned another of the the same lenses and this is ready to switch in next time we need to do a serious realignment of the north cavity. We're leaving it for now as we want to get onto other issues like scatter in the transmission PLL beat path.

I took a reference PLL spectrum attached in the noise budget plot below. There is a new broad 2 kHz bump that we are trying to work out the source of. Previously tara had speculated that this was RFAM (PSL:1311) and Evan had speculated that it was FSS induced noise (PSL:1526). We re optimized the 14.75 MHz RF EOMs polariaziations in both paths reducing North to 80 dBm/sqrtHz and south to 90 dBm/sqrtHz. This didn't seem to be as good as some of my previous reported residual AM, its not clear why North was so difficult to reduce this time; previously North was optimizable down to the noise floor of the PD RF dark noise.  The 2 kHz hump is still there.  It might well be an artifact of the scatter.

Attachment 1: 20171129_135353noiseBudget.pdf  902 kB  Uploaded Tue Dec 5 10:29:08 2017  | Hide | Hide all
20171129_135353noiseBudget.pdf
Attachment 2: 2017-11-29_21.30.17.jpg  1.807 MB  Uploaded Tue Dec 5 10:30:26 2017  | Hide | Hide all
2017-11-29_21.30.17.jpg
Attachment 3: 2017-11-29_21.32.12.jpg  2.411 MB  Uploaded Tue Dec 5 10:30:49 2017  | Hide | Hide all
2017-11-29_21.32.12.jpg
ELOG V3.1.3-